OpenStudy (anonymous):

Which ordered pairs from the replacement set are solutions to the equation? y = -2x + 9 ; Replacement Set {((-1, 11), (0, 8), (1, 7), (2, 5)} A.(-1, 11) B.(0, 8) C.(1, 7) D.(2, 5) *There can be more than one ordered pair. Just double checking my answers.

OpenStudy (mathteacher1729):

Basically you're setting x = -1, x = 0, x = 1, and x = 2 and seeing if you get y = 11, y= 8, y = 7 and y = 5 as the result Example: y = -2x + 9 "the y coordinate is found by taking negative two times the x coordinate and then adding nine to that." let x = -1 y = -2*(-1) + 9 ==> -2 times -1 is +2 y = 2 + 9 add 2 and 9 y = 11 So the ordered pair is (-1, 11) This ordered pair matches the pair in the "replacement set". :)

OpenStudy (anonymous):

Thank You(: What You wrote made a lot of sense.

OpenStudy (mathteacher1729):

Yay! :) Also, I like how you do the smileys backwards. q-;

OpenStudy (anonymous):

Haha and I like your smiley (: lol